# Disentangling Speakers in Multi-Talker Speech Recognition with Speaker-Aware CTC

### v1

@@ Abstract
Multi-talker speech recognition (MTASR) faces unique challenges in disentangling and transcribing overlapping speech. To address these challenges, this paper investigates the role of Connectionist Temporal Classification (CTC) in speaker disentanglement when incorporated with Serialized Output Training (SOT) for MTASR. Our visualization reveals that CTC guides the encoder to represent different speakers in distinct temporal regions of acoustic embeddings. Leveraging this insight, we propose a novel Speaker-Aware CTC (SACTC) training objective, based on the Bayes risk CTC framework. SACTC is a tailored CTC variant for multi-talker scenarios, it explicitly models speaker disentanglement by constraining the encoder to represent different speakers’ tokens at specific time frames. When integrated with SOT, the SOT-SACTC model consistently outperforms standard SOT-CTC across various degrees of speech overlap. Specifically, we observe relative word error rate reductions of 10% overall and 15% on low-overlap speech. This work represents an initial exploration of CTC-based enhancements for MTASR tasks, offering a new perspective on speaker disentanglement in multi-talker speech recognition.  [^1] Code will be open-sourced soon.
@@ I Introduction
Another line of research lies on Serialized Output Training (SOT) [1]. The SOT approach serializes text from different speakers as a single stream, with a special token $\langle sc\rangle$ as a delimiter between speakers. In contrast to the structural priors of BAE models, this approach relies on attention mechanisms in attention encoder-decoder (AED) [13] to disambiguate speakers. This confers an advantage in that it does not require pre-defining speakers and branch numbers, allowing it to handle a variable number of speakers. The superiority of SOT methods has been demonstrated in the M2Met challenges [14, 15], which provided challenging ”in the wild” multi-talker meeting speech. SOT methods have been further enhanced with speaker information [16, 17], time boundary [18, 19], learnable speaker orders [20], and integrated with BAE structures as a hybrid branched SOT model [21].
@@ IV-B Performance of SACTC
We evaluated the proposed SACTC approach using a default risk factor parameter of 15. Initially, we trained a model with the SACTC objective alone. As shown in Table I, SACTC by itself did not outperform the vanilla CTC model (B1 vs. D1). However, when combined with SOT (D1), the model showed significant improvements over vanilla CTC: overall LSM-2mix WER improved from 8.8 to 8.0, and mid-overlap WER from 12.4 to 8.4. This outcome is understandable, as SACTC is designed to enhance MTASR embedding with deterministic speaker disentangling, thus not necessarily improving token-level recognition222 For encoder-only models, there might exist potential trade-offs between these two aspects. . When integrated with SOT, SACTC enhanced low-overlap recognition similar to vanilla CTC, while mitigating performance degradation on more severe overlaps. Experiment E2 also supports this interpretation. For the model trained with SOT+SACTC, AED-only decoding led to performance gains compared to SOT+CTC, particularly improving recognition in high-overlap conditions (13.1→12.3). This suggests that SACTC resulted in embeddings with enhanced speaker discriminability. Based on these findings, we propose that AED-only decoding should be preferred for heavily overlapped scenarios.

### v2

@@ Abstract
Multi-talker speech recognition (MTASR) faces unique challenges in disentangling and transcribing overlapping speech. To address these challenges, this paper investigates the role of Connectionist Temporal Classification (CTC) in speaker disentanglement when incorporated with Serialized Output Training (SOT) for MTASR. Our visualization reveals that CTC guides the encoder to represent different speakers in distinct temporal regions of acoustic embeddings. Leveraging this insight, we propose a novel Speaker-Aware CTC (SACTC) training objective, based on the Bayes risk CTC framework. SACTC is a tailored CTC variant for multi-talker scenarios, it explicitly models speaker disentanglement by constraining the encoder to represent different speakers’ tokens at specific time frames. When integrated with SOT, the SOT-SACTC model consistently outperforms standard SOT-CTC across various degrees of speech overlap. Specifically, we observe relative word error rate reductions of 10% overall and 15% on low-overlap speech. This work represents an initial exploration of CTC-based enhancements for MTASR tasks, offering a new perspective on speaker disentanglement in multi-talker speech recognition. The code is available at https://github.com/kjw11/Speaker-Aware-CTC.
@@ I Introduction
Another line of research lies on Serialized Output Training (SOT) [1]. The SOT approach serializes text from different speakers as a single stream, with a special token $\langle sc\rangle$ as a delimiter between speakers. In contrast to the structural priors of BAE models, this approach relies on attention mechanisms in attention encoder-decoder (AED) [13] to disambiguate speakers. This confers an advantage in that it does not require pre-defining speakers and branch numbers, allowing it to handle a variable number of speakers. The superiority of SOT methods has been demonstrated in the M2Met challenges [14, 15], which provided challenging ”in the wild” multi-talker meeting speech. SOT methods have been further enhanced with speaker information [16, 17], time boundary [18, 19], learnable speaker orders [20], large language models [21], and integrated with BAE structures as a hybrid branched SOT model [22].
@@ IV-B Performance of SACTC
We evaluated the proposed SACTC approach using a default risk factor parameter of 15. Initially, we trained a model with the SACTC objective alone. As shown in Table I, SACTC by itself did not outperform the vanilla CTC model (B1 vs. D1). However, when combined with SOT (D1), the model showed significant improvements over vanilla CTC: overall LSM-2mix WER improved from 8.8 to 8.0, and mid-overlap WER from 12.4 to 8.4. This outcome is understandable, as SACTC is designed to enhance MTASR embedding with deterministic speaker disentangling, thus not necessarily improving token-level recognition111 For encoder-only models, there might exist potential trade-offs between these two aspects. . When integrated with SOT, SACTC enhanced low-overlap recognition similar to vanilla CTC, while mitigating performance degradation on more severe overlaps. Experiment E2 also supports this interpretation. For the model trained with SOT+SACTC, AED-only decoding led to performance gains compared to SOT+CTC, particularly improving recognition in high-overlap conditions (13.1→12.3). This suggests that SACTC resulted in embeddings with enhanced speaker discriminability. Based on these findings, we propose that AED-only decoding should be preferred for heavily overlapped scenarios.
@@
VI Acknowledgements
This work is supported by the HKSARG Research Grants Council’s Theme-based Research Grant Scheme (Project No. T45- 407/19N) and the CUHK Stanley Ho Big Data Decision Research Centre.